OpenStudy (anonymous):

14. For the function rule y = -2x - 3, provide the x-y coordinates of the y-intercept and the x-intercept. Show your work.

OpenStudy (dtan5457):

For the y intercept This is in y=mx+b form y=-2x-3 b= y intercept. For the x intercept, set -2x-3 to zero so -2x-3=0 -2x=3 solve for x.

OpenStudy (anonymous):

y-intercept means x=0 x-intercept means y=0

OpenStudy (anonymous):

how do i solve the x

OpenStudy (anonymous):

can you give me sample how to solve it please

OpenStudy (dtan5457):

-2x=3 divide both sides by -2

OpenStudy (anonymous):

y=mx+b when y=0 => mx=-b when x=0 => y=b

OpenStudy (anonymous):

2x/2=-x 3/2=-1.5 I still dont under stand

OpenStudy (anonymous):

I have to have step by step

OpenStudy (anonymous):

y intercept y=-2x-3 x=0 at y-intercept y=-2(0)-3 y=-3

OpenStudy (anonymous):

x intercept y=-2x-3 y=0 at x intercept 0=-2x-3 add +3 for both sides 3=-2x-3+3 3=-2x divid by -2 for both sides 3/-2=-2/-2x 3/-2=x =>x=-3/2

OpenStudy (anonymous):

which one is which and how do I solve y

OpenStudy (dtan5457):

I'll go over it again.

OpenStudy (dtan5457):

y=-2x-3 y intercept is simply at y=mx+b b is your intercept.

OpenStudy (dtan5457):

For the x intercept -2x-3=0 (set it to zero) -2x=3 divide by -2 x=-1.5

OpenStudy (dtan5457):

understood?

OpenStudy (anonymous):

not really

OpenStudy (dtan5457):

Which part?

OpenStudy (anonymous):

none of it how do i get the inercept y

OpenStudy (dtan5457):

Your original equation is y=-2x-3 right?

OpenStudy (dtan5457):

That linear equation is in the form of y=mx+b

OpenStudy (dtan5457):

When it is in that form, the value of b is the y intercept

OpenStudy (dtan5457):

What is your y intercept..?

OpenStudy (anonymous):

b how do i write it

OpenStudy (dtan5457):

What is the value of "b" in y=-2x-3?

OpenStudy (anonymous):

I dont know alright lets start over from the beginning please cause you telling me to divide 2 to something & I dont know what to divide it too. Can you go step by step but lets do x intercept first then y intercept

OpenStudy (anonymous):

I am sorry it is taken so long to reply my computer keeps acting up

OpenStudy (dtan5457):

Ok so x intercept first.. y=-2x-3 To find the x intercept you must get y as 0... so -2x-3=0 now you need to isolate and solve for x -2x=3 divide by -2 to get rid of the -2 in the x and you get x=-1.5

OpenStudy (anonymous):

i put -2x-3 and it didnt come up with notin

OpenStudy (anonymous):

what do divide he -2 too

OpenStudy (dtan5457):

read my explanation again

OpenStudy (dtan5457):

i can't explain it any better

OpenStudy (anonymous):

Do i write the way you just wrote it

OpenStudy (dtan5457):

sure as long as you understand it

OpenStudy (anonymous):

I dont understand algebra at all

OpenStudy (dtan5457):

i can't really help you unless you tell me which part is actually confusing

OpenStudy (anonymous):

Alright you saying -2x-3=0 how are you getting that

OpenStudy (dtan5457):

The original function was y=-2x-3 to get x intercept, y must equal 0, as every x intercept coordinate is (x,0) so you set it to zero, and solve for x

OpenStudy (anonymous):

how do i solve the x

OpenStudy (dtan5457):

-2x-3=0 put x alone by adding 3 -2x=3 divide by -2 x=-1.5

OpenStudy (anonymous):

So I divide -2x=3 by 2

OpenStudy (dtan5457):

-2

OpenStudy (anonymous):

I just tried to divid that by 2

OpenStudy (dtan5457):

-2x=3 -2 divided by -2=1 or x 3 divided by -2=-1.5 x=-1.25

OpenStudy (anonymous):

got it but from y=-2x-3 how did you get -2x-3=0

OpenStudy (anonymous):

got it now y intercept